# Data-Driven Safety Verification using Barrier Certificates and Matrix Zonotopes

### v1

@@ I Introduction
Ensuring safety is a fundamental requirement in the design and operation of cyber-physical systems (CPSs), particularly in safety-critical applications such as autonomous vehicles, robotics, power grids, and industrial automation [1]. Barrier certificates have emerged as a powerful and widely adopted tool for formal safety verification in dynamical systems [2]. These certificates are Lyapunov-like functions that ensure that system trajectories remain within a predefined safe region and do not reach an unsafe region. A barrier certificate is a real-valued function that is nonpositive over the initial states, positive over the unsafe states, and nonincreasing with transitions. Thus, such a certificate guarantees safety as its zero-level set separates the reachable and unsafe states. Moreover, its zero-sublevel set over-approximates the set of reachable states. By formulating safety conditions as inequalities involving the system’s dynamics, a barrier certificate provides a computationally tractable alternative to reachability analysis. However, traditional approaches [2, 3, 4, 5, 6, 7] assume explicit knowledge of the system model, significantly limiting their applicability in scenarios where the true system dynamics are unknown or uncertain.
@@
To overcome this limitation, recent research has shifted towards data-driven approaches that construct barrier certificates directly from observed system behavior. The authors in [8] introduced a method for synthesizing barrier certificates using the Fundamental Lemma, which allows for the derivation of system properties from input-output data. While this approach removes the need for an explicit system model, it inherently assumes noiseless data, as the Fundamental Lemma relies on exact trajectory observations to infer system behavior. Consequently, its applicability is limited in real-world scenarios, where measurement noise, disturbances, and modeling uncertainties are inevitable. The study in [9, 10] explores data-driven safety verification of unknown discrete-time stochastic systems, presenting a novel approach that formulates the computation of barrier certificates as a robust convex optimization problem. This formulation leverages trajectory-based sampling to construct constraints, ensuring that the derived barrier certificates accurately capture system behavior under uncertainty. The paper in [11] transforms probabilistic barrier certificate constraints into a data-driven optimization framework by defining an ambiguous set of possible transition kernels. To tackle the resulting problem, it leverages sum-of-squares (SOS) optimization [12].
@@
Recent advancements in data-driven reachability analysis have provided a promising alternative to traditional barrier certificate methods. A variety of approaches have emerged to tackle the problem of the unknown model from different perspectives. One approach in [13] introduced two probabilistic methods: one reformulated reachability as a classification problem, while the other applied Monte Carlo sampling to estimate reachable sets. Similarly, [14] proposed a hybrid strategy combining measurement-driven and model-based formal verification techniques. Other works have focused on refining reachability estimates through active learning [15, 16]. A probabilistic framework for general nonlinear systems, based on Christoffel function level sets, was developed in [17]. More recently, neural network-based approaches have gained traction. In [18], feedforward neural networks [19] were employed to approximate reachable sets, with uncertainty quantified via conformal inference [20, 21]. Additionally, kernel density estimation, accelerated by the fast Fourier transform, has been applied to model uncertainties and compute probabilistic reachable sets [22]. A distinct data-driven framework for forward stochastic reachability analysis was introduced in [23]. This method estimates the evolution of the state’s probability density function using trajectory data, which is then used to construct a Gaussian mixture model. Although these probabilistic approaches enhance estimation accuracy with larger datasets, they still lack robust safety guarantees. In particular, rare but safety-critical events remain challenging to capture, limiting the reliability of purely probabilistic reachability methods.
@@
Conversely, researchers have also investigated robust overapproximations of reachability using noise-free data. In [24], interval Taylor-based techniques were applied to systems modeled by differential inclusions, with later extensions incorporating noisy data. Another notable approach, proposed in [25], integrated partial model knowledge with data-driven learning to capture state-dependent uncertainties, relying on the assumption that the unknown dynamics conformed to a known bounding set. Additionally, [26] explored the computation of robust backward reachable sets for unknown linear systems, leveraging noisy data to ensure safety guarantees despite uncertainty. The method proposed in [27, 28] formulates a set of possible system models encapsulated within a matrix zonotope, guaranteeing that every model consistent with the noisy data and prescribed noise bounds is included. However, this approach often results in conservative reachable sets for long horizons. This conservatism can be partially mitigated if there is additional side information available [29]. A key approach to providing infinite horizon safety guarantees is through the use of barrier certificates.
@@
In this paper, we build on the results in [27, 28] to compute a set of models for a linear control system and leverage barrier certificates as a discretization-free approach to provide safety guarantees over an infinite horizon for this given set of models. We use sum-of-squares (SOS) optimization to search for these certificates. To tackle the presence of a set of models, we start by formulating a computationally heavy condition for safety directly from the computed model parameters. Normally, the transition function is a function of the state, input, and noise parameters in some compact set. In the condition we formulate, the transition function is additionally a function of the model parameters that are also in some compact set. This dramatically increases the total number of parameters we need to keep track of, particularly when the barrier certificate is applied over the transition function. The larger memory usage makes it harder to run the optimization problem successfully. Thus, we address this computational challenge of the given condition by stating an equivalent condition that exploits the matrix zonotope representation of the models of the system and scales relatively well with the dimension of the system. We substantiate the benefit of the latter condition by implementing both conditions and investigating their performance for systems with different dimensions. As the models of the system can be made to be less conservative using available side information, our approach can also be integrated with the enhanced model parameters. Moreover, as the computation of matrix zonotope over-approximation of system model parameters has been generalized for polynomial nonlinear control systems, our approach seems promising to be adopted for such systems with appropriate updates. The codes to recreate our findings are publicly available. 111 [Online]. Available: https://github.com/maoumer/Data-Driven-Safety-Verification-using-Barrier-Certificates-and-Matrix-Zonotopes
@@
The remainder of the paper is organized as follows. Section II introduces preliminaries, notation, and problem formulation used in the paper. Section III covers our approach of safety verification via barrier certificates for the data-driven system models while section IV highlights a computation method of finding the barrier certificates via SOS programming. Section V illustrates numerical simulations. Finally, Section VI offers conclusive remarks.
II Preliminaries and Problem Formulation
@@ II-B Set representation
Given $L\in\mathbb{R}^{m\times n}$, the linear transformation of a zonotope $\mathcal{Z}$ is a zonotope $L\mathcal{Z}=\langle Lc_{\mathcal{Z}},LG_{\mathcal{Z}}\rangle$. Given two zonotopes $\mathcal{Z}_{1}=\langle c_{\mathcal{Z}_{1}},G_{\mathcal{Z}_{1}}\rangle\subset\mathbb{R}^{n}$ and $\mathcal{Z}_{2}=\langle c_{\mathcal{Z}_{2}},G_{\mathcal{Z}_{2}}\rangle\subset\mathbb{R}^{n}$, the Minkowski sum and the Cartesian product are respectively computed as [31]:
@@
A matrix zonotope is represented as $\mathcal{M}=\langle C_{\mathcal{M}},G_{\mathcal{M}}\rangle$, and its definition is analogous to that of a zonotope. It is characterized by a center matrix $C_{\mathcal{M}}\in\mathbb{R}^{n\times m}$ and $\gamma_{\mathcal{M}}\in\mathbb{N}$ generator matrices $G_{\mathcal{M}}=\begin{bmatrix}G_{\mathcal{M}_{1}}&\cdots&G_{\mathcal{M}_{\gamma_{\mathcal{M}}}}\end{bmatrix}\in\mathbb{R}^{n\times(m\gamma_{\mathcal{M}})}$ [1, p. 52].
@@
$\begin{bmatrix}Z_{i,j}-M_{i,j}^{min}\\ M_{i,j}^{max}-Z_{i,j}\\ \end{bmatrix}\geq 0.$ (1)
@@ II-D Reachability Analysis
Definition 2 (Exact Reachable Set)
@@
$\forall k\in\{0,...,N{-}1\}\big{\}}.$ (3)
II-E Barrier Certificate
@@
Definition 3 (Barrier Certificate [2])
@@
$\mathcal{B}(x_{k})\leq 0\quad\forall x_{k}\in\mathcal{X}_{0},$ (4)
$\mathcal{B}(x_{k})>0\quad\forall x_{k}\in\mathcal{X}_{u},\text{ and }$ (5)
@@
$\quad\forall u_{k}\in\mathcal{Z}_{u},\forall w_{k}\in\mathcal{Z}_{w}.$ (6)
@@ II-F Problem formulation
where $C_{\mathcal{M}_{w}}\in\mathbb{R}^{n_{x}\times n_{T}}$ and $G_{\mathcal{M}_{w}}\in\mathbb{R}^{n_{x}\times\gamma_{\mathcal{Z}}n_{T}}$. Here, $\mathcal{M}_{w}$ denotes the matrix zonotope resulting from the concatenation of multiple noise zonotopes $\mathcal{Z}_{w}$ [28].
III Data-Driven Safety Verification
This section describes a data-driven safety verification process for an unknown linear system in (2) subject to noise. Barrier certificates [2] use a model of the system to verify safety per condition (3). Due to the presence of noise in the data, no single model can be trusted or made to fit the data precisely to use for the search of a barrier certificate. Instead, we compute a set of models that are consistent with the observed data. Importantly, this set of models is guaranteed to include the true system model, as shown in [28], and it is represented using a matrix zonotope. The following lemma provides a systematic approach to compute this set of models. As we will show later, one can search for barrier certificates for this set of models methodically.
Lemma 1 ([28, Lemma 1])
@@ III Data-Driven Safety Verification
$\mathcal{M}_{\Sigma}=(X_{+}\oplus-\mathcal{M}_{w})D_{-}^{\dagger},$ (7)
@@
$\quad\forall w_{k}\in\mathcal{Z}_{w},\forall A\in[A^{min},A^{max}],\forall B\in[B^{min},B^{max}].$ (8)
While condition (III) is a robust condition for safety verification, using it to find a barrier certificate becomes intractable, particularly as the dimension of the system increases due to the large number of unknown parameters. To address this issue, we use $\mathcal{M}_{\Sigma}$ as follows. $\mathcal{M}_{\Sigma}=\langle C_{\mathcal{M}_{\Sigma}},G_{\mathcal{M}_{\Sigma}}\rangle=\langle C_{\mathcal{M}_{\Sigma}},\mathbb{0}\rangle\oplus\langle\mathbb{0},G_{\mathcal{M}_{\Sigma}}\rangle$. We denote $[A_{c}\ B_{c}]=C_{\mathcal{M}_{\Sigma}}$ and $[A_{G}\ B_{G}]\in\langle\mathbb{0},G_{\mathcal{M}_{\Sigma}}\rangle$. The unknown system dynamics $Ax_{k}+Bu_{k}+w_{k}$ can then be written as $A_{c}x_{k}+B_{c}u_{k}+(A_{G}x_{k}+B_{G}u_{k}+w_{k})$ where $(A_{G}x_{k}+B_{G}u_{k}+w_{k})\in\mathcal{Z}_{d}=(\langle\mathbb{0},G_{\mathcal{M}_{\Sigma}}\rangle)(\mathcal{Z}_{x}\times\mathcal{Z}_{u})+\mathcal{Z}_{w}$. Based on this setup, we provide the following definition for the safety verification of the system using the data-driven bounds provided above.
Definition 4
@@
$\mathcal{B}(x_{k})\leq 0\quad\forall x_{k}\in\mathcal{X}_{0},$ (9)
$\mathcal{B}(x_{k})>0\quad\forall x_{k}\in\mathcal{X}_{u},\text{ and }$ (10)
@@
$\quad\forall u_{k}\in\mathcal{Z}_{u},\forall d_{k}\in\mathcal{Z}_{d}.$ (11)
@@
IV Computation of Safety Certificates
@@
The system dynamics (2) is linear. We say a set $Q\subseteq\mathbb{R}^{n}$ is semi-algebraic if it can be defined with the help of a vector of polynomial inequalities $h(x)$ as $Q=\{x\mid h(x)\geq 0\}$, where the inequalities are element-wise. When the initial set $\mathcal{X}_{0}$ and unsafe set $\mathcal{X}_{u}$ are semi-algebraic [32], conditions (9)-(4) can be cast as a collection of SOS constraints in order to compute a polynomial barrier certificate of a predefined degree. We note that (matrix) zonotopes are semi-algebraic sets. As the conversion of (matrix) zonotopes to their halfspace representation can be exponential in the number of generators [33], we instead convert them to interval matrices as they are scalable and computationally convenient due to their ease of representation as polynomial inequalities.
Assumption 3
@@
$-\mathcal{B}(x)-\lambda_{0}^{T}(x)g_{0}(x),$ (12)
$\mathcal{B}(x)-\epsilon-\lambda_{u}^{T}(x)g_{u}(x),$ (13)
$\mathcal{B}(x)-\mathcal{B}(A_{c}x+B_{c}u+d)-\lambda^{T}(x,u,d)g(x,u,d),$ (14)
@@
V Numerical simulations
In this section, we demonstrate the application of our approach in a case study. The simulations were conducted on a Windows 11 device equipped with an AMD Ryzen 9 4900HS Mobile Processor and 16 GB of RAM. We used CORA v2020 [34] in MATLAB for the matrix zonotope related computations and transferred the resulting model data to Julia where we used TSSOS [35] to implement the SOS optimization problem. We first obtain $\mathcal{M}_{\Sigma}$ and attempt to search for a barrier certificate using conditions (9), (10), and (III). For state $x\in\mathbb{R}^{n_{x}},u\in\mathbb{R}^{n_{u}}$, the total number of parameters according to these conditions is $n_{x}\times n_{x}+n_{x}+n_{x}\times n_{u}+n_{u}+n_{x}=n_{x}^{2}+n_{x}n_{u}+2n_{x}+n_{u}$. This does not scale favorably with respect to the state dimension $n_{x}$ and input dimension $n_{u}$. With each additional degree of the polynomial template for the barrier certificate, the memory needed scales exponentially over these parameters. Thus, it was not computationally tractable for our case study beyond a two dimensional system. Note that condition (14) in Lemma 2 is replaced with

### v2

@@ 1. Introduction
Ensuring safety is a fundamental requirement in the design and operation of cyber-physical systems (CPSs), particularly in safety-critical applications such as autonomous vehicles, robotics, power grids, and industrial automation [Alt10]. Barrier certificates have emerged as a powerful and widely adopted tool for formal safety verification in dynamical systems [PJ04]. These certificates are Lyapunov-like functions that ensure that system trajectories remain within a predefined safe region and do not reach an unsafe region. A barrier certificate is a real-valued function that is nonpositive over the initial states, positive over the unsafe states, and nonincreasing with transitions. Thus, such a certificate guarantees safety as its zero-level set separates the reachable and unsafe states. Moreover, its zero-sublevel set over-approximates the set of reachable states. By formulating safety conditions as inequalities involving the system’s dynamics, a barrier certificate provides a computationally tractable alternative to reachability analysis. However, traditional approaches [PJ04, ACE+19, JSZ20, WHE18, PJP07, PJP04] assume explicit knowledge of the system model, significantly limiting their applicability in scenarios where the true system dynamics are unknown or uncertain.
@@
To overcome this limitation, recent research has shifted towards data-driven approaches that construct barrier certificates directly from observed system behavior. The authors in [BvH25] introduced a method for synthesizing barrier certificates using the Fundamental Lemma, which allows for the derivation of system properties from input-output data. While this approach removes the need for an explicit system model, it inherently assumes noiseless data, as the Fundamental Lemma relies on exact trajectory observations to infer system behavior. Consequently, its applicability is limited in real-world scenarios, where measurement noise, disturbances, and modeling uncertainties are inevitable. The study in [SZ22, SLSZ24] explores data-driven safety verification of unknown discrete-time stochastic systems, presenting a novel approach that formulates the computation of barrier certificates as a robust convex optimization problem. This formulation leverages trajectory-based sampling to construct constraints, ensuring that the derived barrier certificates accurately capture system behavior under uncertainty. The paper in [SZS24] transforms probabilistic barrier certificate constraints into a data-driven optimization framework by defining an ambiguous set of possible transition kernels. To tackle the resulting problem, it leverages sum-of-squares (SOS) optimization [Par03].
@@
Recent advancements in data-driven reachability analysis have provided a promising alternative to traditional barrier certificate methods. A variety of approaches have emerged to tackle the problem of the unknown model from different perspectives. One approach in [DA20] introduced two probabilistic methods: one reformulated reachability as a classification problem, while the other applied Monte Carlo sampling to estimate reachable sets. Similarly, [HAVdH15] proposed a hybrid strategy combining measurement-driven and model-based formal verification techniques. Other works have focused on refining reachability estimates through active learning [CRDD18, CDDR20]. A probabilistic framework for general nonlinear systems, based on Christoffel function level sets, was developed in [DYEGA21]. More recently, neural network-based approaches have gained traction. In [HQLD23], feedforward neural networks [Saz06] were employed to approximate reachable sets, with uncertainty quantified via conformal inference [VGS05, LW14]. Additionally, kernel density estimation, accelerated by the fast Fourier transform, has been applied to model uncertainties and compute probabilistic reachable sets [WMC25]. A distinct data-driven framework for forward stochastic reachability analysis was introduced in [CPH24]. This method estimates the evolution of the state’s probability density function using trajectory data, which is then used to construct a Gaussian mixture model. Although these probabilistic approaches enhance estimation accuracy with larger datasets, they still lack robust safety guarantees. In particular, rare but safety-critical events remain challenging to capture, limiting the reliability of purely probabilistic reachability methods.
@@
Conversely, researchers have also investigated robust overapproximations of reachability using noise-free data. In [DVG+22], interval Taylor-based techniques were applied to systems modeled by differential inclusions, with later extensions incorporating noisy data. Another notable approach, proposed in [FAZ+19], integrated partial model knowledge with data-driven learning to capture state-dependent uncertainties, relying on the assumption that the unknown dynamics conformed to a known bounding set. Additionally, [AL23] explored the computation of robust backward reachable sets for unknown linear systems, leveraging noisy data to ensure safety guarantees despite uncertainty. The method proposed in [AKAJ21, AKAJ23] formulates a set of possible system models encapsulated within a matrix zonotope, guaranteeing that every model consistent with the noisy data and prescribed noise bounds is included. However, this approach often results in conservative reachable sets for long horizons. This conservatism can be partially mitigated if there is additional side information available [AJS+22]. A key approach to providing infinite horizon safety guarantees is through the use of barrier certificates.
@@
In this paper, we build on the results in [AKAJ21, AKAJ23] to compute a set of models for a linear control system and leverage barrier certificates as a discretization-free approach to provide safety guarantees over an infinite horizon for this given set of models. We use sum-of-squares (SOS) optimization to search for these certificates. To tackle the presence of a set of models, we start by formulating a computationally heavy condition for safety directly from the computed model parameters. Normally, the transition function is a function of the state, input, and noise parameters in some compact set. In the condition we formulate, the transition function is additionally a function of the model parameters that are also in some compact set. This dramatically increases the total number of parameters we need to keep track of, particularly when the barrier certificate is applied over the transition function. The larger memory usage makes it harder to run the optimization problem successfully. Thus, we address this computational challenge of the given condition by stating an equivalent condition that exploits the matrix zonotope representation of the models of the system and scales relatively well with the dimension of the system. We substantiate the benefit of the latter condition by implementing both conditions and investigating their performance for systems with different dimensions. As the models of the system can be made to be less conservative using available side information, our approach can also be integrated with the enhanced model parameters. Moreover, as the computation of matrix zonotope over-approximation of system model parameters has been generalized for polynomial nonlinear control systems, our approach seems promising to be adopted for such systems with appropriate updates. The codes to recreate our findings are publicly available. 111 [Online]. Available: https://github.com/maoumer/Data-Driven-Safety-Verification-using-Barrier-Certificates-and-Matrix-Zonotopes
@@
The remainder of the paper is organized as follows. Section 2 introduces preliminaries, notation, and problem formulation used in the paper. Section 3 covers our approach of safety verification via barrier certificates for the data-driven system models while section 4 highlights a computation method of finding the barrier certificates via SOS programming. Section 5 illustrates numerical simulations. Finally, Section 6 offers conclusive remarks.
2. Preliminaries and Problem Formulation
@@ 2.2. Set representation
Given $L\in\mathbb{R}^{m\times n}$, the linear transformation of a zonotope $\mathcal{Z}$ is a zonotope $L\mathcal{Z}=\langle Lc_{\mathcal{Z}},LG_{\mathcal{Z}}\rangle$. Given two zonotopes $\mathcal{Z}_{1}=\langle c_{\mathcal{Z}_{1}},G_{\mathcal{Z}_{1}}\rangle\subset\mathbb{R}^{n}$ and $\mathcal{Z}_{2}=\langle c_{\mathcal{Z}_{2}},G_{\mathcal{Z}_{2}}\rangle\subset\mathbb{R}^{n}$, the Minkowski sum and the Cartesian product are respectively computed as [FFA+24]:
@@
A matrix zonotope is represented as $\mathcal{M}=\langle C_{\mathcal{M}},G_{\mathcal{M}}\rangle$, and its definition is analogous to that of a zonotope. It is characterized by a center matrix $C_{\mathcal{M}}\in\mathbb{R}^{n\times m}$ and $\gamma_{\mathcal{M}}\in\mathbb{N}$ generator matrices $G_{\mathcal{M}}=\begin{bmatrix}G_{\mathcal{M}_{1}}&\cdots&G_{\mathcal{M}_{\gamma_{\mathcal{M}}}}\end{bmatrix}\in\mathbb{R}^{n\times(m\gamma_{\mathcal{M}})}$ [Alt10, p. 52].
@@
(1) $\begin{bmatrix}Z_{i,j}-M_{i,j}^{min}\\ M_{i,j}^{max}-Z_{i,j}\\ \end{bmatrix}\geq 0.$
@@ 2.4. Reachability Analysis
Definition 2 (Exact Reachable Set).
@@
(3) $\forall k\in\{0,...,N{-}1\}\big{\}}.$
2.5. Barrier Certificate
@@
Definition 3 (Barrier Certificate [PJ04]).
@@
(4) $\mathcal{B}(x_{k})\leq 0\quad\forall x_{k}\in\mathcal{X}_{0},$
(5) $\mathcal{B}(x_{k})>0\quad\forall x_{k}\in\mathcal{X}_{u},\text{ and }$
@@
(6) $\quad\forall u_{k}\in\mathcal{Z}_{u},\forall w_{k}\in\mathcal{Z}_{w}.$
@@ 2.6. Problem formulation
where $C_{\mathcal{M}_{w}}\in\mathbb{R}^{n_{x}\times n_{T}}$ and $G_{\mathcal{M}_{w}}\in\mathbb{R}^{n_{x}\times\gamma_{\mathcal{Z}}n_{T}}$. Here, $\mathcal{M}_{w}$ denotes the matrix zonotope resulting from the concatenation of multiple noise zonotopes $\mathcal{Z}_{w}$ [AKAJ23].
3. Data-Driven Safety Verification
This section describes a data-driven safety verification process for an unknown linear system in (2) subject to noise. Barrier certificates [PJ04] use a model of the system to verify safety per condition (3). Due to the presence of noise in the data, no single model can be trusted or made to fit the data precisely to use for the search of a barrier certificate. Instead, we compute a set of models that are consistent with the observed data. Importantly, this set of models is guaranteed to include the true system model, as shown in [AKAJ23], and it is represented using a matrix zonotope. The following lemma provides a systematic approach to compute this set of models. As we will show later, one can search for barrier certificates for this set of models methodically.
Lemma 1 ([AKAJ23, Lemma 1]).
@@ 3. Data-Driven Safety Verification
(7) $\mathcal{M}_{\Sigma}=(X_{+}\oplus-\mathcal{M}_{w})D_{-}^{\dagger},$
@@
(8) $\quad\forall w_{k}\in\mathcal{Z}_{w},\forall A\in[A^{min},A^{max}],\forall B\in[B^{min},B^{max}].$
While condition (3) is a robust condition for safety verification, using it to find a barrier certificate becomes intractable, particularly as the dimension of the system increases due to the large number of unknown parameters. To address this issue, we use $\mathcal{M}_{\Sigma}$ as follows. $\mathcal{M}_{\Sigma}=\langle C_{\mathcal{M}_{\Sigma}},G_{\mathcal{M}_{\Sigma}}\rangle=\langle C_{\mathcal{M}_{\Sigma}},\mathbb{0}\rangle\oplus\langle\mathbb{0},G_{\mathcal{M}_{\Sigma}}\rangle$. We denote $[A_{c}\ B_{c}]=C_{\mathcal{M}_{\Sigma}}$ and $[A_{G}\ B_{G}]\in\langle\mathbb{0},G_{\mathcal{M}_{\Sigma}}\rangle$. The unknown system dynamics $Ax_{k}+Bu_{k}+w_{k}$ can then be written as $A_{c}x_{k}+B_{c}u_{k}+(A_{G}x_{k}+B_{G}u_{k}+w_{k})$ where $(A_{G}x_{k}+B_{G}u_{k}+w_{k})\in\mathcal{Z}_{d}=(\langle\mathbb{0},G_{\mathcal{M}_{\Sigma}}\rangle)(\mathcal{Z}_{x}\times\mathcal{Z}_{u})+\mathcal{Z}_{w}$. Based on this setup, we provide the following definition for the safety verification of the system using the data-driven bounds provided above.
Definition 4.
@@
(9) $\mathcal{B}(x_{k})\leq 0\quad\forall x_{k}\in\mathcal{X}_{0},$
(10) $\mathcal{B}(x_{k})>0\quad\forall x_{k}\in\mathcal{X}_{u},\text{ and }$
@@
(11) $\quad\forall u_{k}\in\mathcal{Z}_{u},\forall d_{k}\in\mathcal{Z}_{d}.$
@@
4. Computation of Safety Certificates
@@
The system dynamics (2) is linear. We say a set $Q\subseteq\mathbb{R}^{n}$ is semi-algebraic if it can be defined with the help of a vector of polynomial inequalities $h(x)$ as $Q=\{x\mid h(x)\geq 0\}$, where the inequalities are element-wise. When the initial set $\mathcal{X}_{0}$ and unsafe set $\mathcal{X}_{u}$ are semi-algebraic [BCR13], conditions (9)-(4) can be cast as a collection of SOS constraints in order to compute a polynomial barrier certificate of a predefined degree. We note that (matrix) zonotopes are semi-algebraic sets. As the conversion of (matrix) zonotopes to their halfspace representation can be exponential in the number of generators [ASB10], we instead convert them to interval matrices as they are scalable and computationally convenient due to their ease of representation as polynomial inequalities.
Assumption 3.
@@
(12) $-\mathcal{B}(x)-\lambda_{0}^{T}(x)g_{0}(x),$
(13) $\mathcal{B}(x)-\epsilon-\lambda_{u}^{T}(x)g_{u}(x),$
(14) $\mathcal{B}(x)-\mathcal{B}(A_{c}x+B_{c}u+d)-\lambda^{T}(x,u,d)g(x,u,d),$
@@
5. Numerical simulations
In this section, we demonstrate the application of our approach in a case study. The simulations were conducted on a Windows 11 device equipped with an AMD Ryzen 9 4900HS Mobile Processor and 16 GB of RAM. We used CORA v2020 [Alt15] in MATLAB for the matrix zonotope related computations and transferred the resulting model data to Julia where we used TSSOS [WML21] to implement the SOS optimization problem. We first obtain $\mathcal{M}_{\Sigma}$ and attempt to search for a barrier certificate using conditions (9), (10), and (3). For state $x\in\mathbb{R}^{n_{x}},u\in\mathbb{R}^{n_{u}}$, the total number of parameters according to these conditions is $n_{x}\times n_{x}+n_{x}+n_{x}\times n_{u}+n_{u}+n_{x}=n_{x}^{2}+n_{x}n_{u}+2n_{x}+n_{u}$. This does not scale favorably with respect to the state dimension $n_{x}$ and input dimension $n_{u}$. With each additional degree of the polynomial template for the barrier certificate, the memory needed scales exponentially over these parameters. Thus, it was not computationally tractable for our case study beyond a two dimensional system. Note that condition (14) in Lemma 2 is replaced with